OpenStudy (anonymous):

I am really struggling with this one! It involves fractions. I need to solve it algebraically. 1/4=3-2x-1/x+2. the last part of this equation is a fraction.

OpenStudy (anonymous):

is it 1 1 ----- = 3-2x - -------- ??? 4 x-2

OpenStudy (anonymous):

not quite 2x-1 over x+2

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

1 2x - 1 ----- = 3 - -------- 4 x - 2 1 3(x-2) - (2x -1) ----- = -------------- 4 x-2 1 3x- 6 - 2x +1 ----- = -------------- 4 x-2 1 x - 5 ----- = -------- 4 x-2

OpenStudy (anonymous):

cross-multiplying 1(x-2) = 4(x-5) x-2 = 4x - 20 -2 + 20 = 4x - x 18 = 3x 18/3 = x 6 = x

OpenStudy (anonymous):

if u r satisfied with the solution, pls click GOOD ANSWER button for me... (:-}))

OpenStudy (anonymous):

\[\frac{1}{4}=3-\frac{2x-1}{x+2}\]?

OpenStudy (anonymous):

oh yeah, i made a mistake then........

OpenStudy (anonymous):

\[\frac{1}{4}-3=-\frac{2x-1}{x+2}\] \[-\frac{11}{4}=-\frac{2x-1}{x+2}\] \[\frac{11}{4}=\frac{2x-1}{x+2}\] \[11(x+2)=4(2x-1)\] \[11x+22=8x-4\] \[3x=-26\] \[x=-\frac{26}{3}\]

OpenStudy (anonymous):

1 2x - 1 ----- = 3 - -------- 4 x + 2 1 3(x+2) - (2x -1) ----- = -------------- 4 x+2 1 3x+ 6 - 2x +1 ----- = -------------- 4 x+2 1 x + 7 ----- = -------- 4 x+2

OpenStudy (anonymous):

cross-multiplying 1(x+2) = 4(x+7) x+2 = 4x + 28 2 - 28 = 4x - x -26 = 3x -26/3 = x

OpenStudy (anonymous):

Okay guys and girls. i am not sure where the good answer button is. Sorry this is my first time.

OpenStudy (anonymous):

harkirat is correct, but no need to add fractions if you want to just cross multiply

OpenStudy (anonymous):

the GOOD ANSWER button is on the extreme right of the bar above this answer, in which my name is written..........

OpenStudy (anonymous):

I would rather cross multiply

OpenStudy (anonymous):

r u using IE browser??

OpenStudy (anonymous):

nevermind it allowed me to vote. Thanks

OpenStudy (anonymous):

u r welcome :))